In a loading van the driver pushed a 15kg crate for 5 meters up the ramp. She had to push with a force of 132N and the vertical height gained was 1.3m. What was the efficiency of pushing this crate up the ramp onto the back of the van?
1
Expert's answer
2021-10-28T08:52:19-0400

"F = 132N"

"s = 5m"

"m= 15kg"

"h= 1.3m"

"g=9.8\frac{m}{s^2}"

"\eta= \frac{W_{output}}{W_{input}}*100 \%"

"W_{output }= mgh = 15*9.8*1.3= 191.1J"

"W_{input}=Fs=132*5=660J"

"\eta= \frac{191.1}{660}*100\%\approx 29\%"


"\text{Answer: }\eta = 29\%"